OpenStudy (anonymous):

expand in simplest form. log base 5 of ab

jimthompson5910 (jim_thompson5910):

use the rule \[\Large \log_{b}(xy) = \log_{b}(x)+\log_{b}(y)\] to get this \[\Large \log_{5}(ab) = \log_{5}(a)+\log_{5}(b)\]

OpenStudy (anonymous):

what is the change of base formula?

jimthompson5910 (jim_thompson5910):

the change of base formula is \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\] you could also state it using natural logs too \[\Large \log_{b}(x) = \frac{\ln(x)}{\ln(b)}\] since the logs on the right side can be of any allowed base (as long as they are the same base)

OpenStudy (anonymous):

so in this case log base 15 of 17 would 15 be b?

jimthompson5910 (jim_thompson5910):

correct, b = 15 and x = 17 \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\] \[\Large \log_{15}(17) = \frac{\log(17)}{\log(15)}\]

OpenStudy (anonymous):

i tried that and it said its not the answer:/

jimthompson5910 (jim_thompson5910):

sounds like they want you to use a calculator to evaluate \(\Large \frac{\log(17)}{\log(15)}\)

OpenStudy (anonymous):

i dont have a fancy one

jimthompson5910 (jim_thompson5910):

use google then type in "log(17)/log(15)" without quotes into google and it will show a calculator with the result inside

jimthompson5910 (jim_thompson5910):

tell me what result you get so I can confirm

OpenStudy (anonymous):

1.04

jimthompson5910 (jim_thompson5910):

you should have gotten 1.04621891533

jimthompson5910 (jim_thompson5910):

if you round that to two decimal places, then you would go from 1.04621891533 to 1.05

OpenStudy (anonymous):

so would log base 13 of 15 be 0.947?

jimthompson5910 (jim_thompson5910):

you have the fraction mixed up

jimthompson5910 (jim_thompson5910):

log base 13 of 15 = log(15)/log(13) = ???

OpenStudy (anonymous):

ohh okay i got it. i did the next three. number 11. the hydrogen concentration [H+] in a certain cleaning compound is [H+]=3.7*10^-11. use the formula pH= -log[H+] to find the pH of the cleaning compound

jimthompson5910 (jim_thompson5910):

you just need to evaluate -1*log(3.7 * 10^(-11))

OpenStudy (anonymous):

10.432?

OpenStudy (anonymous):

what is the properties of logarithms? i need it for log base 3 of 27- 2 log base 3 of 3

jimthompson5910 (jim_thompson5910):

27 = 3^3

jimthompson5910 (jim_thompson5910):

log(27) = log(3^3) log(27) = 3*log(3)

jimthompson5910 (jim_thompson5910):

I'm using the rule that log(x^y) = y*log(x)

OpenStudy (anonymous):

so 1.431

jimthompson5910 (jim_thompson5910):

hmm let me think

jimthompson5910 (jim_thompson5910):

\[\Large \log_{3}(27) - 2\log_{3}(3)\] \[\Large \log_{3}(3^3) - 2\log_{3}(3)\] \[\Large 3\log_{3}(3) - 2\log_{3}(3)\] \[\Large 3*1 - 2*1\] \[\Large 3 - 2\] \[\Large 1\] --------------------------------------- So, \[\Large \log_{3}(27) - 2\log_{3}(3)=1\]

OpenStudy (anonymous):

okay i did 12-13... 14. write the following logarithm as a quotient of two common logarithms. Do not simplify. log base 2 of 15

jimthompson5910 (jim_thompson5910):

use the change of base formula and tell me what you get

jimthompson5910 (jim_thompson5910):

they don't want you to use a calculator

OpenStudy (anonymous):

log15/log2

jimthompson5910 (jim_thompson5910):

yep or log(15)/log(2) and that's all they want

OpenStudy (anonymous):

expand the following. log base 8 of sqrt r+6/ s^2t^1/4

jimthompson5910 (jim_thompson5910):

can you draw that out?

OpenStudy (anonymous):